OpenStudy (anonymous):

Flip one coin and then roll one die. What is the probability of getting a tail on the coin and a number less than 5 on the die?

OpenStudy (anonymous):

I'm thinking this is 0 because the two events are mutually exclusive.

OpenStudy (anonymous):

The answer must be a fraction might I add.

OpenStudy (kropot72):

The probability of getting a tail on the coin is 1/2. The probability of getting a number less than 5 on the die is 4/6. The two events are independent so the combined probability is their product: \[\frac{1}{2}\times \frac{4}{6}=?\]

OpenStudy (kropot72):

@tinyneko Do you understand?

OpenStudy (anonymous):

THANK YOU! And yes I do.

OpenStudy (kropot72):

You're welcome :)